Solve,
x + y = 6
y = 2x - 3
I'm completely lost
Daren

Answer by ankor@dixie-net.com(22740)

Let's see if we can find you.
:
Solve,
x + y = 6
y = 2x - 3
:
Note that the 2nd equation tell us y = (2x-3)
so replace y with (2x-3) in the 1st equation
x + (2x-3) = 6
3x - 3 = 6
add 3 to both sides
3x = 6 + 3
3x = 9
x = 9/3
x = 3
:
Find y using y = 2x - 3, replace x with 3, and you have
y = 2(3) - 3
y = 6 - 3
y = 3
:
:
We can check our solutions of x=3, y=3 in the 1st equation x + y = 6
3 + 3 = 6. You can check in the 2nd equation